OpenStudy (anonymous):

I DO NOT WANT YOU TO GIVE ME THE ANSWER! Sol’s Sweater Emporium is having a sale on sweaters. The following prices are posted in the window. Number of Sweaters Total Cost 3 $72 5 $120 10 $240 Which of the following best describes the unit cost per sweater? A. Each sweater costs $18. <--- B. Each sweater costs $22. C. Each sweater costs $24. D. Each sweater costs $28.

OpenStudy (xguardians):

You want me to walk you through this one :)?

OpenStudy (xguardians):

They are all ratios. $72 (Cost) / 3 (Amount Purchased) = ???? $120 (Cost) / 5 (Amount Purchased) = ????? $240 (Cost) / 10 (Amount Purchased) = ????? Solve any of them for the answer.

OpenStudy (xguardians):

@CrazyCountryGirl

OpenStudy (anonymous):

take the average of both sides and divide the right one by the left one!

OpenStudy (anonymous):

xgames is easier though

OpenStudy (anonymous):

1) Add the total cost 2)Add the number of sweaters 3)Now divide the total cost by number of sweater to get the unit cost

OpenStudy (anonymous):

@xGuardians Yes please. :)

OpenStudy (anonymous):

$72/3=24 $120/5=24 $240/10=24 Yeah that makes more sense. :) Thanks!